OpenStudy (ayee_ciera):

A candle is 17 in. tall after burning for 3 hours. After 5 hours, it is 15 in. tall. Write a linear equation to model the relationship between height h of the candle and time t. Predict how tall the candle will be after 8 hours.

OpenStudy (campbell_st):

so can you find the slope the 2 points are (3, 17) and(5, 15)

OpenStudy (campbell_st):

once you know the slope use the slope point formula to get the equation of the line.

OpenStudy (ayee_ciera):

the slope is -1

OpenStudy (campbell_st):

great so you have a point (3, 17) and slope of -1 so the equation is y = -x + b substitute y = 17 and x = 3 to find b

OpenStudy (campbell_st):

you many need to write it as h = -t + b and b is the y intercept or initial height

OpenStudy (ayee_ciera):

@campbell_st i dont get it

OpenStudy (campbell_st):

ok... so if you use the point slope formula \[y - y_{1} = m(x - x_{1})\] using m = -1 and (3, 17) y - 17 = -1(x - 3) y - 17 = -x + 3 now you should be able to find the equation

OpenStudy (campbell_st):

if you use the slope intercept form of a line y = mx + b or using the slope m = -1 y = -x + b you have a point and slope 17 = -1(3) + b so b = 20 so the equation is y = -x + 20 so just substitute x = 8 to find the height after 8 hours

OpenStudy (ayee_ciera):

since the -x is a negative should the 8 be negative as well